I must solve the equation using the add/subtract method
3x+9y=9
-x+5y=-3

Answer by stanbon(75887)

I must solve the equation using the add/subtract method
3x+9y=9
-x+5y=-3
------
Multiply thru the lower equation by 3:
3x+9y = 9
-3x+15y = -9
-----
Add the two equations and solve for "y":
25y = 0
y = 0
----
Substitute into -x+5y = -3 and solve for "x":
-x+0 = -3
x = 3
------
Cheers,
Stan H.